MOA-II Obsevation Log

Return to index
Previous day Next day

DATE: 2007/06/19
OBSERVER: Kamiya Nagaya Paul
CONDITION: clear -> fog
COMMENT: D2359(60sec) is OK. Shutter fixed. camera mounted. We executed telescope analysis. chip2 bias value suddenly high ~2000ADU. Resistance between the camera and the telescope is about 9kohm, but CCD noise is not increased.
Compressor gas pressure: 1.8 MPa
Expected Observation Time: START 18:20 END 07:18


TARGETRUNTOTAL RUNTOTAL FIELDTOTAL EXPOSURE (sec)
GB7806-78061160
LMC-000
SMC-000
ALERT-000
GRB-000
TOTAL---1160 (00h01m00s)


GB FIELD STATISTICS

Field1Field2Field3Field4Field5Field6Field7Field8Field9Field10Field11Field12
000010000000
Field13Field14Field15Field16Field17Field18Field19Field20Field21Field22Field23
00000000000


LMC FIELD STATISTICS

Field1Field2Field3Field4Field5Field6Field7Field8Field9Field10Field11Field12
000000000000
Field13Field14Field15Field16
0000


SMC FIELD STATISTICS

Field1Field2Field3Field4Field5Field6Field7Field8Field9Field10Field11Field12
000000000000


SPECIAL TARGET NAMEBR>
CategoryTargetNrun


CHECK SHEET

TimeStatusWeather
18:30waitingclear
19:00waitingclear
20:00waitingclear,fog
21:00waitingclear
22:00waitingclear
23:00waitingclear
0:00waitingclear
1:00waitingclear
2:00waitingclear
3:00waitingclear, tpoint
4:00waitingclear
5:00waitingclear
6:00waitingfog
7:00waitingfog, dawn


CCD TEMPERATURE Previous day Next day




VACUUM Previous day Next day




TELESCOPE TEMPERATURE Previous day Next day




Return to index

Previous day Next day